Estimate the number of gallons of gasoline consumed by the total of all automobile drivers in the U.S., per year. (There are about 3 x 108 people in the U.S. Assume that half of them have cars, that they drive an average of 12,000 miles per year, and that their cars get an average of 20 miles per gallon of gasoline.)


1
Expert's answer
2021-10-02T14:55:35-0400

"\text{cars}=\frac{\text{people}}2=1.5\cdot 10^8,"

"\text{total miles}=\text{mpy}\cdot \text{cars}=18\cdot 10^{11},"

"\text{gallons}=\frac{\text{gpy}}{\text{mpg}}=9\cdot 10^9."